The second hand on a clock is 6 inches long. How far does the tip of the hand move in 40 seconds?
Answer by stanbon(48522)

The second hand on a clock is 6 inches long. How far does the tip of the hand move in 40 seconds?
------
Circumference = 2*Pi*r
(40/60)(2 pi r) = (2/3)(2 pi 6)
=8pi inches
-------------
Cheers,
Stan H.